In a new developement, 50 one-and two-bedroom condominiums were sold. Each one-bedroom condominium sold for $120 thousand and each two bedroom condominium sold for $150 thousand. If sales totaled $7050 thousand, how many of each type of unit was sold?
Answer by Alan3354(69443)

In a new developement, 50 one-and two-bedroom condominiums were sold. Each one-bedroom condominium sold for $120 thousand and each two bedroom condominium sold for $150 thousand. If sales totaled $7050 thousand, how many of each type of unit was sold?
---------------------
n = one bedroom
t = 2 bedroom
n+t = 50 --> t = 50-n
120n + 150t = 7050
Sub for t
120n + 150*(50-n) = 7050
12n + 750 - 15n = 705
-3n = -45
n = 15
----------
t = 35